OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (anonymous):

@ANAS23 comment and ill give you the medal

OpenStudy (anonymous):

seems complicated but not hard. So the facts u given is that he is 10 years older and tim played football 3 years ago making coach twice his age. i prob cant give an equation. but i can get u an answer. the answer is that tim was tim was 10 when he started. when he started tim was 10 and the coach was 20 thats double. than 3 years later tim is 13 and coach is 23.

OpenStudy (anonymous):

1) C = coach; T = Tim C = T + 10; C-3 = 2(T-3), so C = 2T - 6 + 3 = 2t - 3 2t - 3 = t + 10 t = 13, Tim's age now, so coach is 23; three years ago they were 10 & 20

OpenStudy (anonymous):

my medal

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Let's assign: Coach Johnson's age : J Tim's age: T Now according to our data we have: Coach Johnson is 10 years older than Tim That means: J = T+10 Also, three years ago, Coach Johnson was twice his age.. That means: J-3 = 2*(T-3) Let's solve this set of equations using substitution: J-3 = (T+10)-3 = 2T -6 => T=7+6 => T = 13 therefore Tim's age is 13 years old J=T+10 => J=23 therefore coach Johnson's age is 23 years old.. This is correct since they have 10 years difference and three years ago Tim was 10 and Johnson was 20, so tthe coach had twice Tim's age. Solved! :D

OpenStudy (anonymous):

mine is easier to understand

OpenStudy (anonymous):

@Fox is right.

OpenStudy (anonymous):

true^ @Fox375 comment on my next one and ill give you a meda;. it doesnt have to be correct ou deserve a medal :)

OpenStudy (anonymous):

medal*

OpenStudy (nubeer):

lol why so desperate for medals.?

OpenStudy (anonymous):

i need points fast

OpenStudy (anonymous):

im not, they are x) idunno why lolz

OpenStudy (anonymous):

My same question @nubeer lol

OpenStudy (nubeer):

lol i was asking them.. medals dont help to increase your smart score.

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

fiesty... XD

OpenStudy (anonymous):

whatev

OpenStudy (anonymous):

@Lyndsey_Coach_xox thaks